Three ice skaters meet at the center of a rink and each stands at rest facing the center, within arm's reach of the other two. On a signal, each skater pushes himself away from the other two across the frictionless ice. After the push, skater A with mass mA = 70.0 kg

 moves in the negative y-direction at 2.00 m/s and skater B with mass mB = 85.0 kg

 moves in the negative x-direction at 3.50 m/s. Find the x- and y-components of the 75.0 kg skater C's velocity (in m/s) after the push.



vCx=__ m/s

vCy=___ m/s


1
Expert's answer
2021-10-03T13:24:44-0400

Three ice skaters meet at the center of a rink and we can consider that they are at rest before they start to move thus "v_{\text{skater A}_i} = v_{\text{skater B}_i} = v_{\text{skater C}_i} = 0\,m/s", in consequence "\vec{p}_{\text{skater C}_i}=\vec{p}_{\text{skater B}_i}=\vec{p}_{\text{skater A}_i}=\vec{p_{i}}=\vec{0}" .


Since momentum is conserved, this means that "\vec{p_i}=\vec{p_f}=\vec{p}_{\text{skater A}_f}+\vec{p}_{\text{skater B}_f}+\vec{p}_{\text{skater C}_f}=\vec{0}". Now, we need to define the final conditions for the three skaters considering that A and B are moving along the negative y and x-axis, respectively:


"\vec{v}_{\text{skater A}_f} =[0\frac{m}s,-2\frac{m}s]; m_{\text{A}} =70.0\,kg
\\ \vec{v}_{\text{skater B}_f} =[-3.5\frac{m}s,0\frac{m}s]; m_{\text{B}} =85.0\,kg
\\ \vec{v}_{\text{skater C}_f} =[v_{c_x},v_{c_y}]; m_{\text{C}} =75.0\,kg"


Considering this we substitute for "\vec{p_f}= \vec{0}" with the provided information when they started to move:


"\\ m_{A}\vec{v}_{\text{skater A}_f} +m_{B}\vec{v}_{\text{skater B}_f} + m_{C}\vec{v}_{\text{skater C}_f} =\vec{0}
\\ (70\,kg) [0\frac{m}s,-2\frac{m}s]+(85\,kg)[-3.5\frac{m}s,0\frac{m}s]+ (75\,kg)[v_{c_x},v_{c_y}] =\vec{0}
\\ [0,-140](\frac{kgm}s)+[-297.5,0](\frac{kgm}s)+ (75\,kg)[v_{c_x},v_{c_y}] =\vec{0}"


Following we can find "[v_{c_x},v_{c_y}]" as


"[v_{c_x},v_{c_y}]= -\frac{(\frac{kgm}s)}{(75\,kg)}( [0,-140]+[-297.5,0])
\\ [v_{c_x},v_{c_y}]= -\frac{1}{(75)} [-297.5,-140](\frac{m}s)
\\ [v_{c_x},v_{c_y}]= [+\frac{119}{30},+\frac{28}{15}](\frac{m}s)\approxeq[+3.966,+1.866](\frac{m}s)"



In conclusion, vCx= 3.966 m/s and vCy= 1.866 m/s. The skater C traves through the xy quadrant at vC= 4.384 m/s.



Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.